OpenStudy (anonymous):

Consider the following system of equation: x + y − z =2, x + 2y + z =3, x + y + (a^2−5)z =a, where a ∈ R. Find all values of a for which the resulting linear system has (a) no solution, (b) a unique solution, (c) infinitely many solutions.

OpenStudy (anonymous):

OMG it's Tori Black... LOL

OpenStudy (ybarrap):

For a unique solution, the determinant of these equation needs to be non zero. For an infinite number of solutions, we must reduce to 2 equations with 3 unknowns. For no solutions, one or more of these planes would be parallel.

OpenStudy (ybarrap):

do you know about determinants?

OpenStudy (anonymous):

no

OpenStudy (ybarrap):

Ok then. For a unique solution, all rows should be independent. Choose an "a" that makes this true.

OpenStudy (anonymous):

i solved the system and got a unique solution of \[\frac{ 1 }{ 2 }(1\pm \sqrt(17) \]

OpenStudy (anonymous):

so how do i find no solutions and infinitely many

OpenStudy (ybarrap):

You have 4 unknowns, x,y, z and a but only 3 equations. There are no unique solutions until "a" is fixed. For certain "a" any of these three possibilities are possible (unique, infinite,none).

OpenStudy (anonymous):

yeah i found "a" for a unique solution

OpenStudy (anonymous):

could you solve it?

OpenStudy (ybarrap):

yes

OpenStudy (ybarrap):

If a=2, then last row would be equal to the 1st row, right? What doe that mean?

OpenStudy (ybarrap):

What does that imply about the independence of these rows? Are they dependent or independent? When a=2.

OpenStudy (tkhunny):

...or -2.

OpenStudy (ybarrap):

correct.

OpenStudy (tkhunny):

On the other hand, a = 2 is not as catastrophic as a = -2.

OpenStudy (ybarrap):

With a = 2, you find that there is infinitely many solutions. So how do we find one "a" that has no solution. We need to choose an a that makes the last equation "parallel" to any of the other two equations.

OpenStudy (anonymous):

so any value other than the unique solution value and a=+-2 should ake it parallel?

OpenStudy (anonymous):

when a=2 there are NOT infinitely many solutions

OpenStudy (ybarrap):

My line of thinking is this. Each of the coefficients of this equation represents a vector that is perpendicular to the planes each equation represents. If two equations have the same perpendicular vector then they are also parallel to each other. Does this make sense?

OpenStudy (ybarrap):

When we used a=2, we found the last row equal to the first row, right? But what if a = -2?

OpenStudy (anonymous):

yes sorry made an error

OpenStudy (ybarrap):

With a=-2 we have two planes that are parallel, but not the same plane. This means...?

OpenStudy (ybarrap):

This means we found a plane parallel to another but doesn't intersect it. That is, no solution.

OpenStudy (anonymous):

what is the value of "a" when there are no solutions?

OpenStudy (ybarrap):

a=-2 is one possibility that I see

OpenStudy (anonymous):

but there could be many values? how would i right that?

OpenStudy (ybarrap):

a=2 is infinite solution

OpenStudy (anonymous):

something like: all values excluding a=-1 and a=2 will give no solutions?

OpenStudy (ybarrap):

Look's like a=-1 would give a unique solution

OpenStudy (anonymous):

and values that would give no solutions?

OpenStudy (ybarrap):

a=-2 would give no solutions. I don't see another possibility for no solutions

OpenStudy (ybarrap):

What I'm looking for is an a that makes the last row a linear combination of one of the 1st two but the constants are not equal

OpenStudy (tkhunny):

\[\left[\begin{matrix}1 & 0 & 0 & \dfrac{a+5}{a+2} \\ 0 & 1 & 0 & \dfrac{a}{a+2}? \\ 0 & 0 & 1 & \dfrac{1}{a+2}\end{matrix}\right]\] Only a = -2 is that big a problem.

OpenStudy (anonymous):

thanks @tkhunny

OpenStudy (tkhunny):

Oops, I forgot to delete a question mark and I forgot to say that was Reduced Row Echelon form. Darn that I.E. I had to switch to Firefox just to see what I was typing!

OpenStudy (ybarrap):

lol. In any case, without some tools in linear algebra this is probably the best I can do to describe. I hope this helped